OpenStudy (magbak):

Solve for x: 2x2 - 4x - 14 = 0 (1 point)

OpenStudy (luigi0210):

Factor out a 2 first

OpenStudy (magbak):

\[This is what I got x= (2-4\sqrt{2})/ 2\]

OpenStudy (luigi0210):

How did you get that?

OpenStudy (magbak):

I simplified a equation that a= 1 b= -2 c= -7

OpenStudy (luigi0210):

Wait, you solved it using the quadratic formula?

OpenStudy (magbak):

Yes I think so.

OpenStudy (anonymous):

You can factor out the 2, and that will help a little, or you actually can just use the numbers given. Personally, I would factor the 2 out and come down to:\[2(x ^{2} - 2x - 7) = 0\]Now you have: a = 1, b = -2, and c = -7 You are correctly using the quadratic formula:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]When you are done, don't forget to see if you can further simplify.

OpenStudy (anonymous):

Also, don't forget that there is a pos/neg sign in the quadratic formula. So, your answer was close, but you need to include the pos also. And again, don't forget to simplify further.

OpenStudy (magbak):

Yes but when I solved for this I got \[x= (2-4\sqrt{2}\div) 2\]

OpenStudy (magbak):

Yes their were two equation the one I gave you and then one with a positive 4

OpenStudy (anonymous):

\[x = \frac{ -(-2) \pm \sqrt{(-2)^{2} - (4)(1)(-7)} }{ (2)(1) }\]\[x = \frac{ 2 \pm \sqrt{4 + 28} }{ 2 }\]I'll let you simplify it from here, but somewhere, you will divide the numerator and denominator by "2".

OpenStudy (magbak):

ok when I take the four out side can I add the 2 and the 4 or not.

OpenStudy (anonymous):

\[x = \frac{ 2 \pm \sqrt{32} }{ 2 } = \frac{ 2 \pm 4\sqrt{2} }{ 2 } = 1 \pm 2\sqrt{2}\]

OpenStudy (magbak):

Ok so that is the final answer.

OpenStudy (anonymous):

Yes. Because the numerator and the denominator can be divided by "2", we can cancel out the "2" from both. Yes, that's the final answer. But more importantly, those are the steps.

OpenStudy (magbak):

I appreciate you even more each time.

OpenStudy (anonymous):

So, at some time, you might want to review how the +- stayed in all the calculations. You're very good to work with! Nice person! @magbak